# Arg-LLaDA: Argument Summarization via Large Language Diffusion Models and Sufficiency-Aware Refinement

### v1

@@ Abstract
Hao Liα Yizheng Sunα Viktor Schlegel Kailai Yangα Riza Batista-Navarroα Goran Nenadicα α The University of Manchester, UK β Imperial Global Singapore, Imperial College London, Singapore Correspondence: hao.li-2@manchester.ac.uk
@@ 3.2 Diffusion Generation Framework
This objective encourages the model to generate contextually appropriate and semantically grounded content.
Inference-Time Sufficiency Guided Masking.
A key strength of the diffusion framework lies in its capacity for selective revision across denoising steps. Rather than masking tokens at random during inference, we introduce a sufficiency-guided masking strategy that targets semantically weak or unsupported spans. While it is theoretically possible to incorporate sufficiency-guided masking during training, we deliberately apply it only at inference time for both practical and technical reasons. First, introducing sufficiency feedback into training would require dense span-level sufficiency annotations or reliable pseudo-labels—resources that are typically scarce and noisy. Second, training-time masking is expected to be diverse and general-purpose, enabling the model to learn broadly applicable denoising capabilities. Conditioning masking decisions too early on sufficiency scores may introduce optimization instability and reduce generalizability. By contrast, applying sufficiency-guided masking only during inference allows us to retain a modular and interpretable framework, where the sufficiency signal can flexibly adapt to different evaluation conditions without requiring retraining. This separation of training and inference responsibilities follows prior work on diffusion generation (Gonzalez et al., 2023) and proves especially effective for iterative refinement tasks such as argument summarization.
To operationalize this, we integrate a sufficiency diagnosis module (Section 3.3) that estimates token-level sufficiency scores , reflecting the degree to which each token is justified by the input arguments. These scores define a corruption distribution used to select the next resampling mask :
where introduces sampling diversity and balances guided corruption and exploration. At each step, we sample the top- proportion of tokens according to for regeneration. This approach enables fine-grained control over the generation trajectory and improves factual sufficiency without sacrificing fluency.
Given an input comprising claims and their supporting evidence, we aim to generate a coherent abstractive summary by denoising a corrupted summary through a learned reverse process. At each timestep , the model samples an intermediate state from the previous state using the reverse transition kernel . This process is repeated until convergence:
The initial corrupted summary is fully masked (i.e., all tokens replaced by a [MASK] symbol), ensuring that generation is driven entirely by learned knowledge and the input context .
3.3 Sufficiency Diagnosis Module.
To enable targeted refinement, we introduce a sufficiency diagnosis module that detects unsupported or redundant spans in the generated summary. A summary span is deemed sufficient if it is adequately grounded in the input claims and evidence. We consider two complementary approaches to assess sufficiency: an LLM-based reasoning prompt and a supervised classification model.
Chain-of-Thought Prompting.
@@
Inspired by recent advances in reasoning with large language models, we construct few-shot prompts to elicit explicit sufficiency judgments. Given a candidate span and its supporting arguments , the model is prompted to reason about whether is (i) fully supported, (ii) insufficient or unverifiable, or (iii) redundant with other spans. This approach improves transparency and provides interpretable rationales for refinement decisions.
Supervised Sufficiency Classifier
To enable span-level refinement, we develop a supervised sufficiency classifier inspired by factuality checking frameworks in text generation. This classifier plays a key role in guiding iterative refinement by identifying spans that lack sufficient grounding in the source arguments.
Formally, for a summary span , associated claim , and supporting evidence , the classifier predicts a binary label indicating sufficiency:
where is a RoBERTa-based encoder and denotes input concatenation with segment indicators. The model is trained to minimize binary cross-entropy loss over annotated examples.
@@ 3.4 Training and Implementation Details
We adopt a two-stage training strategy for Arg-LLaDA, comprising diffusion-based denoising and sufficiency classification.
@@
We initialize the diffusion model with the LLaDA checkpoint (Nie et al., 2025), pretrained on large-scale web corpora using masked diffusion objectives. Its architecture and noise scheduling are retained. For fine-tuning, we supervise the model on domain-specific argument summarization data. Given input tuples of claims and evidence with reference summaries, we apply stochastic masking (30% ratio). The model is then optimized to denoise the masked summary using the diffusion objective (Section 3.2).
@@
We train a RoBERTa-base encoder (Liu et al., 2019) to predict span-level sufficiency labels. Each training instance comprises a summary span , an associated claim , and supporting evidence . Labels are automatically constructed through controlled perturbation, including hallucinated, contradictory, and unsupported variants. The classifier is trained with binary cross-entropy loss and used during inference to guide masking decisions.
@@ 5 Results and Analysis
As shown in Table 3 (right), removing the sufficiency diagnosis module (“No Diagnosis”) results in a significant drop across all evaluation metrics—ROUGE-L drops from to , BLEURT from to , and BERTScore from to . This indicates that without explicit sufficiency filtering, the generation becomes more verbose and includes unsupported content, thereby reducing both faithfulness and conciseness.

### v3

@@ Abstract
Hao Li1 Yizheng Sun1 Viktor Schlegel1,2 Kailai Yang1 Riza Batista-Navarro1 Goran Nenadic1 1 The University of Manchester, UK 2 Imperial Global Singapore, Imperial College London, Singapore Correspondence: kailai.yang@manchester.ac.uk
@@ 3.2 Diffusion Generation Framework
This objective encourages the model to generate contextually appropriate and semantically grounded content. Importantly, masking during training remains random to ensure broad denoising capability, while targeted refinement is deferred to inference.
3.3 Inference-Time Sufficiency Guided Masking
In line with recent advances in diffusion-based text generation, Arg-LLaDA adopts an inference-time selective masking strategy to refine summaries across iterative denoising steps. Instead of relying solely on random corruption, our framework integrates two complementary signals—token-level confidence and span-level sufficiency. This combined design allows the model to address both locally uncertain expressions and confidently generated but unsupported statements, thereby achieving more faithful and concise summaries.
Token-level low-confidence remasking. At each inference step, the model produces probability distributions over candidate tokens. We compute token confidence as the maximum probability assigned to any candidate token. Tokens with confidence below a threshold are deemed unreliable, and the lowest- proportion of such tokens are selected for resampling. This mechanism ensures that local corrections are directed toward genuinely uncertain regions while avoiding unnecessary revisions of fluent and stable tokens. Following prior work on diffusion-based generation (Nie et al., 2025), we apply confidence-based remasking only at inference, not training. This separation ensures that the diffusion backbone learns a broad, task-agnostic denoising capability, while inference-time confidence signals specialize revision to context-specific uncertainties.
Sufficiency-guided span masking. Low-confidence detection alone cannot address confidently incorrect content or omissions. To mitigate this, we integrate an external sufficiency module , which evaluates whether candidate summary spans are adequately supported by the input claims and evidence. Spans are segmented primarily at the sentence level, aligned with punctuation and discourse boundaries. For compound sentences, we further split into clause-level units to obtain semantically coherent argumentative statements. Each span is assigned a sufficiency score or label, and tokens within spans flagged as insufficient are projected into the corruption distribution.
Combined corruption distribution. Both signals are unified into a single masking distribution:
where is the token-level confidence, the span-level sufficiency assigned to the token’s enclosing span, and a balancing coefficient. At each denoising step, we sample tokens for resampling according to this distribution.
3.4 Sufficiency Diagnosis Module
To implement the global component of inference-time masking, we introduce a sufficiency diagnosis module that evaluates whether generated spans are adequately grounded in the input arguments. Here, a summary span is defined at the sentence level (or clause-level segmentation for long sentences) to ensure coherent semantic units. A span is considered sufficient if it is consistent with the provided claims and evidence while avoiding redundancy with other spans. This module provides span-level sufficiency scores that complement token-level confidence, enabling the system to revise both uncertain tokens and confidently produced but unsupported statements. We explore two complementary approaches.
@@ 3.4 Sufficiency Diagnosis Module
Chain-of-Thought Prompting. We construct few-shot prompts for large language models to explicitly reason about sufficiency. Given a candidate span and its supporting arguments , the model is prompted to determine whether is fully supported, insufficient, or redundant. This prompting strategy provides interpretable rationales and flexible adaptation across domains, though it may introduce variability depending on model scale and prompting quality.
Supervised Sufficiency Classifier. In parallel, we develop a RoBERTa-based classifier to provide stable sufficiency signals. For each span , claim , and evidence , the classifier predicts a binary sufficiency label:
where is the encoder and denotes concatenation with segment indicators. Training data are constructed via controlled perturbations (hallucinated, contradictory, or unsupported variants) to ensure coverage of common insufficiency types. During inference, the classifier provides span-level sufficiency scores that are integrated into the corruption distribution, complementing low-confidence token masking.
@@ 3.5 Training and Implementation Details
We adopt a two-stage training strategy for Arg-LLaDA, comprising diffusion-based denoising and sufficiency classification. Importantly, sufficiency guidance is introduced only at inference time: training uses diverse random masking to teach general denoising skills, while inference selectively leverages sufficiency signals for targeted refinement. This separation avoids overfitting to noisy sufficiency pseudo-labels and maintains generalization.
@@
We initialize the diffusion model with the LLaDA checkpoint (Nie et al., 2025), pretrained on large-scale web corpora with masked diffusion objectives. Its architecture and noise scheduling are retained. For fine-tuning, we supervise the model on domain-specific argument summarization data. Given input tuples of claims and evidence with reference summaries, we apply stochastic masking (30% ratio). The model is then optimized to denoise the masked summary using the diffusion objective (Section 3.2).
@@
We train a RoBERTa-base encoder (Liu et al., 2019) to predict span-level sufficiency labels. Each training instance consists of a sentence-level span , its associated claim , and supporting evidence . Labels are automatically generated through controlled perturbations, including hallucinations, contradictions, and omissions. The classifier is trained with binary cross-entropy loss and applied only during inference to guide selective span masking.
@@ 5 Results and Analysis
As shown in Table 4, removing the sufficiency diagnosis module (“No Diagnosis”) results in a significant drop across all evaluation metrics—ROUGE-L drops from to , BLEURT from to , and BERTScore from to . This indicates that without explicit sufficiency filtering, the generation becomes more verbose and includes unsupported content, thereby reducing both faithfulness and conciseness.